# Gradient-Sign Masking for Task Vector Transport Across Pre-Trained Models

### v1

@@ 4.3 Limited Data Regime
In Sec. 4.1, we have assumed access to the full target dataset $\mathcal{D}$ to compute the gradient $\bm{g}$} at the zero-shot target model $\theta_{B}$. In practice, one of the main motivations for task vector transport is the few-shot or limited data regime. If we had access to the entire dataset, we could directly fine-tune $\theta_{B}$ to obtain $\theta_{B}^{ft}$, making task vector transfer efficient but sub-optimal.
When only a small number of samples is available, we estimate the gradient signs using a subset of labeled examples. Let $\mathcal{D}_{s}\subset\mathcal{D}$ denote a small subset of $N$ samples. For each parameter coordinate $i$, we compute the sign of the gradient via majority voting across these samples:
$\hat{s}_{i}=\mathrm{sign}\left({\sum}_{(x_{n},y_{n})\in\mathcal{D}_{s}}\mathrm{sign}\left(\nabla_{\theta}\ell(f_{\theta_{B}}(x_{n}),y_{n})\right)\right).$ (9)
@@
Let $p_{i}=\Pr[\mathrm{sign}(\nabla_{\theta}\ell(f_{\theta_{B}}(x),y))=\mathrm{sign}(\tau_{B,i})]$ denote the probability that a single-sample gradient aligns with the oracle task vector $\tau_{B}$. Then, under mild independence assumptions, $p_{i}>1/2$, and the majority vote estimator satisfies the following:
$\Pr\left[\mathrm{sign}(g_{i})=\mathrm{sign}(\tau_{B,i})\right]\geq 1-\exp\left(-2N(p_{i}-1/2)^{2}\right),$ (10)
which shows that the estimated sign concentrates around the true gradient direction as the number of samples $N$ grows.
We provide a proof of this lemma in Appendix A, which makes use of Hoeffding’s inequality (Hoeffding, 1963). In practice, even a few samples provide a robust estimate of the true gradient direction. Each gradient acts as a “vote” for the correct sign, and majority voting filters out noisy or conflicting directions. This guarantees that, with high probability, the masked task vector $\delta_{B}^{mask}$ points in a descent direction, preserving the first-order loss reduction property of the full-data update. As shown in Sec. 5.2, this approach is robust to small sample sizes, making it particularly attractive when direct fine-tuning of $\theta_{B}$ is expensive or prone to overfitting.
@@
7:Build gradient-sign mask: $m_{i}\leftarrow\mathbbm{1}\{\mathrm{sign}(\tau_{A,i})=\hat{s}_{i}\}$
@@
9:return $\theta_{B}^{trans}\leftarrow\theta_{B}-\delta^{A}$ $\triangleright$ Updated target model
@@ 5 Experimental Results
Implementation details. For the Vision Settings, we consider CLIP ViT-B/16 and ViT-L/14 Vision Transformers (Radford et al., 2021), implemented in Open-CLIP (Cherti et al., 2023). The original pre-trained weights are denoted $\theta_{A}$ and the target model weights $\theta_{B}$. For ViT-B/16, $\theta_{A}$ was pre-trained on Datacomp XL (s13b, b90k) and $\theta_{B}$ on LAION-2B (s34b, b88k). For ViT-L/14, $\theta_{A}$ was pre-trained on Datacomp XL (s13b, b90k) and $\theta_{B}$ on LAION-2B (s32b, b82k). For the Language Settings, we investigated different Text-To-Text Transfer Transformer (T5) (Raffel et al., 2020) models in the base configuration. As $\theta_{A}$, we used T5v1.1, pre-trained on the C4 (Raffel et al., 2020) dataset without any supervised training. For $\theta_{B}$, we used FLAN-T5 (Chung et al., 2024), pre-trained and instruction-tuned on several datasets, including GSM8K (Cobbe et al., 2021), AQUA-RAT (Ling et al., 2017), and LAMBADA (Paperno et al., 2016). Task vectors were obtained following the fine-tuning protocol of Ilharco et al. (2023): $2000$ iterations, batch size $128$, learning rate $1\times 10^{-5}$, cosine annealing with $200$ warm-up steps, AdamW optimizer (Loshchilov & Hutter, 2019), weight decay $0.1$. The text encoder backbone was kept frozen following Cherti et al. (2023).
Baselines. We evaluate our method against several baselines. As a lower bound, we consider the zero-shot target model ($\theta_{B}$ zero-shot), i.e., the base model without any fine-tuning. As an upper bound, we report $\theta_{B}+\delta^{\star}$, obtained by adding the source task vector $\tau_{A}$ masked with the signs of the true task vector $\tau_{B}$ to the target model. We also include the performance of the fully fine-tuned target model ($\theta_{B}$ fine-tune) and the naive task arithmetic transport ($\theta_{B}+\tau_{A}$). In addition, we compare against TransFusion (Rinaldi et al., 2025), which transports task vectors across transformer-based models via permutation alignment. Finally, we report the accuracy of a target model fine-tuned with the same number of randomly sampled examples per class $|D_{s}|$ used by our approach.
Supervision Budget $\mathcal{D}_{s}$. In all experiments, the subset $\mathcal{D}_{s}$ is drawn from the full downstream fine-tuning dataset $\mathcal{D}$ and constitutes only a fraction of its size. Throughout the tables, $|\mathcal{D}_{s}^{c}|$ indicates the number of examples per class used to estimate gradient signs for the target model $\theta_{B}$. The corresponding proportions of $\mathcal{D}$ used to form $\mathcal{D}_{s}$ are provided in Appendix C.
Table 1: Cross-dataset performance comparison between ViT-B/16 and ViT-L/14 models.
@@ 5.1 Transport Experiments
Transport in the Vision Setting. Tab. 1 summarizes the results of task vector transport across CLIP ViT-B/16 and CLIP ViT-L/14 architectures, averaged over multiple random seeds that determine the composition of the sampled $\mathcal{D}_{s}$ (standard deviations are reported in Sec. B.3). Our GradFix , denoted by $\theta_{B}-\delta^{A}$, yields a consistent improvement over naive task vector addition ($\theta_{B}+\tau_{A}$) even when using a single sample per class to approximate true gradient signs. Notably, the naive addition performs nearly at the level of zero-shot initialization and fails to transfer meaningful task knowledge. This confirms that GradFix effectively suppresses misaligned components of $\tau_{A}$, preventing negative transfer due to pre-training mismatch.
To further evaluate our approach, we compare it against few-shot fine-tuning of $\theta_{B}$, denoted as $\theta_{B}^{opt}$, using the same limited target samples. GradFix achieves better performance, on both ViT-B/16 and ViT-L/14, while exhibiting smaller variance across seeds with respect to few-shot fine-tuning. Moreover, as the $\mathcal{D}_{s}$ size increases, our method continues to provide stable gains, whereas $\theta_{B}^{opt}$ suffers from fluctuations and instability due to the composition of the supervision dataset. These results demonstrate that our approach ensures consistent and reliable task vector transport, remaining stable across different subsets $\mathcal{D}_{s}$. Importantly, this robustness is achieved with a single forward–backward pass to obtain the mask $\bm{m}$}, highlighting the efficiency and simplicity of the proposed method.
Table 2: Cross-dataset performance of T5 models on different NLP tasks
@@
$\theta_{B}-\delta^{A}$ $50$ $68.06$ $49.68$ $54.25$ $60.50$ $59.89$ $58.48$
Transport in the Language Setting. Tab. 2 reports results on task vector transport across T5 models evaluated on closed-vocabulary text classification benchmarks. While direct addition of $\tau_{A}$ to $\theta_{B}$ fails to transfer knowledge effectively, our method substantially closes the gap toward full fine-tuning, confirming its ability to identify and retain task-relevant directions. Notably, the relative improvement over naive transfer is even larger than in the vision setting, underscoring the robustness of our approach in domains where task transfer is especially challenging. This demonstrates that the benefits of GradFix are not confined to vision, and that a single forward–backward pass suffices to enable reliable and efficient task vector transport also in the language domain.
@@ 5.2 Masking Strategies
Table 3: Performance of $\theta_{B}$ with oracle or estimated gradient signs under different mask strategies: agreement retains matching signs, force agreement aligns all signs, and random assigns signs uniformly. Results averaged over seeds with $|\mathcal{D}_{s}^{c}|=1$ on CLIP ViT-B/16.
@@
$\theta_{B}+\delta^{\star}$ force agreement $97.95$ $93.51$ $95.94$ $96.60$ $80.59$ $92.92$
agreement $95.06$ $87.06$ $82.92$ $92.04$ $71.44$ $85.71$
$\theta_{B}-\delta^{A}$ force agreement $61.32$ $70.10$ $60.91$ $70.52$ $58.05$ $64.18$
agreement $61.94$ $70.05$ $60.89$ $71.07$ $58.32$ $64.45$
$\theta_{B}-\delta^{A}$ random $49.49$ $67.97$ $48.41$ $50.54$ $56.06$ $54.50$
@@
To analyze the effect of different mask sign strategies on the transport of the task vector $\tau_{A}$, we compare our mask construction method, denoted agreement, with two alternatives: force agreement and random masks. Tab. 3 reports results using 1 sample per class on ViT-B/16, averaged across multiple random seeds. As reference points, we include the zero-shot base model ($\theta_{B}$) and the fully fine-tuned model ($\theta_{B}$ fine-tune), providing lower and upper bounds on performance.
@@
For both $\delta^{A}$ and $\delta^{\star}$, the mask $\bm{m}$} determines which directions of $\tau_{A}$ are retained Eq. 5. In agreement, $\bm{m}$} retains only the coordinates whose signs match those of the reference as in Eq. 4. In force agreement, all signs of $\tau_{A}$ are aligned with the signs of the gradient $\bm{s}$}, obtaining the mask:
$m_{i}^{fa}=\mathrm{sign}(\tau_{A,i})\cdot\hat{s}_{i},$ (11)
@@
flipping any entries that disagree with the reference. When the oracle $\tau_{B}$ is used as the reference, force agreement generally outperforms agreement, as fully leveraging the true task direction maximizes transfer. In contrast, using few-shot gradient-based estimates from $\theta_{B}$, agreement performs slightly better than force agreement. This is consistent with the fact that the gradient-based estimate is noisy. Forcing all directions can propagate errors, while keeping only agreeing entries provides a more reliable mask. Finally, we evaluate random masks, where the signs are randomly sampled from a uniform distribution $\mathcal{U}\{-1,+1\}$. The random masks yield performance close to the zero-shot baseline, confirming that unstructured perturbations provide no meaningful guidance.

### v3

@@ 4.3 Limited Data Regime
In Sec. 4.1, we have assumed access to the full target dataset $\mathcal{D}$ to compute the gradient $\bm{g}$} at the zero-shot target model $\theta_{B}$. In practice, one of the main motivations for task vector transport is the few-shot or limited data regime. If we had access to the entire dataset, we could directly fine-tune $\theta_{B}$ to obtain $\theta_{B}^{ft}$, making task vector transfer unnecessary in that setting.
When only a small number of samples is available, we estimate the anti-gradient signs using a subset of labeled examples. Let $\mathcal{D}_{s}\subset\mathcal{D}$ denote a small subset of $N$ samples. For each parameter coordinate $i$, we compute the sign of the anti-gradient via majority voting across these samples:
$\hat{s}_{i}=\operatorname{sign}\left(-{\sum}_{(x_{n},y_{n})\in\mathcal{D}_{s}}\operatorname{sign}\left(\nabla_{\theta}\ell\left(f_{\theta_{B}}(x_{n}),y_{n}\right)\right)\right).$ (9)
@@
Let $p_{i}=\Pr[\operatorname{sign}(\nabla_{\theta}\ell(f_{\theta_{B}}(x),y))=\operatorname{sign}(g_{i})]$ denote the probability that a single-sample gradient sign matches the true gradient sign at coordinate $i$. Then, under mild independence assumptions and for $p_{i}>1/2$, the majority-vote estimator satisfies:
$\Pr\left[\hat{s}_{i}=\operatorname{sign}(-g_{i})\right]\geq 1-\exp\left(-2N{(p_{i}-1/2)}^{2}\right),$ (10)
which shows that the estimated sign concentrates around the true anti-gradient direction as the number of samples $N$ grows.
We provide a proof of this lemma in Appendix A, which uses Hoeffding’s inequality (hoeffding1963probability). In practice, even a few samples provide a robust estimate of the true descent direction. Each gradient acts as a vote for the correct sign, and majority voting filters out noisy or conflicting directions. This implies that, with high probability, the masked task vector $\delta^{A}$ points in a descent direction, preserving the first-order loss reduction behavior of the full-data update. As shown in Sec. 5.3, this approach is robust to small sample sizes, making it particularly attractive when direct fine-tuning of $\theta_{B}$ is expensive or prone to overfitting. Compared to mean-based aggregation, majority voting is less sensitive to outlier magnitudes because it depends only on sign frequency and provides a more stable transfer (Sec. 5.4).
@@
7:Build gradient-sign mask: $m_{i}\leftarrow\mathbbm{1}\{\operatorname{sign}(\tau_{A,i})=\hat{s}_{i}\}$
@@
9:return $\theta_{B}^{trans}\leftarrow\theta_{B}+\delta^{A}$ $\triangleright$ Updated target model
@@ 5 Experimental Results
Implementation details. For the vision settings, we consider CLIP ViT-B/16 and ViT-L/14 Vision Transformers (radford2021learning), implemented in Open-CLIP (cherti2023reproducible). The source pre-trained weights are denoted $\theta_{A}$ and the target pre-trained weights are denoted $\theta_{B}$. For ViT-B/16, $\theta_{A}$ was pre-trained on Datacomp XL (s13b, b90k) and $\theta_{B}$ on LAION-2B (s34b, b88k). For ViT-L/14, $\theta_{A}$ was pre-trained on Datacomp XL (s13b, b90k) and $\theta_{B}$ on LAION-2B (s32b, b82k). For the language settings, we use T5-base variants (raffel2020exploring). As $\theta_{A}$, we use T5v1.1, pre-trained on C4 (raffel2020exploring) without supervised training. As $\theta_{B}$, we use FLAN-T5 (chung2024scaling), pre-trained and instruction-tuned on several datasets, including GSM8K (cobbe2021training), AQUA-RAT (ling2017program), and LAMBADA (paperno2016lambadadatasetwordprediction). Task vectors were obtained following the fine-tuning protocol of ilharco2023task: $2000$ iterations, batch size $128$, learning rate $1\times 10^{-5}$, cosine annealing with $200$ warm-up steps, AdamW optimizer (loshchilov2018decoupled), weight decay $0.1$. The text encoder backbone was kept frozen following cherti2023reproducible. For fairness, all compared transport methods use the same source/target checkpoints and the same supervision budget $\mathcal{D}_{s}$, and results are aggregated across random seeds. The $\alpha$ is chosen through a validation set following standard model merging pipelines, more details in Appendix E.
Baselines. We evaluate our method against several baselines. As a lower bound, we consider the zero-shot target model ($\theta_{B}$ zero-shot), i.e., the base model without fine-tuning. As an upper bound, we report $\theta_{B}+\delta^{\star}$, obtained by adding the source task vector $\tau_{A}$ masked with the signs of the true task vector $\tau_{B}$ to the target model. We also include the performance of the fully fine-tuned target model ($\theta_{B}$ fine-tune) and the naive Task Arithmetic transport ($\theta_{B}+\tau_{A}$). In addition, we compare against TransFusion (rinaldi2025update), which transports task vectors across transformer-based models via permutation alignment. Finally, we report the performance of a target model fine-tuned with the same number of randomly sampled examples per class $|\mathcal{D}_{s}^{c}|$ used by our approach.
Supervision budget $\mathcal{D}_{s}$. In all experiments, the subset $\mathcal{D}_{s}$ is drawn from the full downstream fine-tuning dataset $\mathcal{D}$ and constitutes only a fraction of its size. Throughout the tables, $|\mathcal{D}_{s}^{c}|$ indicates the number of examples per class used to estimate anti-gradient signs for the target model $\theta_{B}$. The corresponding proportions of $\mathcal{D}$ used to form $\mathcal{D}_{s}$ are provided in Appendix D. Additionally, we include an ablation study in Sec. C.2 analyzing the impact of different subset selection heuristics on anti-gradient sign estimation.
Table 1: Cross-dataset performance comparison for ViT-B/16 and ViT-L/14 models.
@@ 5.1 Transport Experiments
Transport in the vision setting. Table 1 summarizes the results of task vector transport across CLIP ViT-B/16 and CLIP ViT-L/14 architectures, averaged over multiple random seeds that determine the composition of the sampled $\mathcal{D}_{s}$ (standard deviations are reported in Sec. B.1). Our GradFix, denoted by $\theta_{B}+\delta^{A}$, yields a consistent improvement over naive task vector addition ($\theta_{B}+\tau_{A}$) even when using a single sample per class to approximate true anti-gradient signs. Notably, the naive addition performs nearly at the level of zero-shot initialization and fails to transfer meaningful task knowledge. This confirms that GradFix effectively suppresses misaligned components of $\tau_{A}$, preventing negative transfer due to pre-training mismatch.
To further evaluate our approach, we compare it against few-shot fine-tuning of $\theta_{B}$, denoted as $\theta_{B}^{opt}$, using the same limited target samples and number of training steps (see Appendix F for a detailed computational cost analysis). This comparison isolates the effect of update construction: both methods observe the same supervision budget, but $\theta_{B}^{opt}$ relies on iterative parameter optimization, whereas GradFix applies a single masked transport step. GradFix achieves better performance on both ViT-B/16 and ViT-L/14, while exhibiting smaller variance across seeds with respect to few-shot fine-tuning (Tabs. 5 and 6). Moreover, as the $\mathcal{D}_{s}$ size increases, our method continues to provide stable gains, whereas $\theta_{B}^{opt}$ shows larger fluctuations due to subset composition. These trends suggest that GradFix is not only stronger on average, but also more reliable under realistic data-sampling variability. In practical low-shot settings, this matters because adaptation quality should remain stable even when the selected subset is not carefully curated. Overall, gradient-sign masking yields a more data-efficient and predictable adaptation mechanism, where robustness is achieved with a single forward-backward pass to obtain the mask $\bm{m}$}.
Table 2: Cross-dataset performance of T5 models on different NLP tasks.
@@
$\theta_{B}+\delta^{A}$ $50$ $68.06$ $49.68$ $54.25$ $60.50$ $59.89$ $58.48$
Transport in the language setting. Table 2 reports results on task vector transport across T5 models evaluated on closed-vocabulary text classification benchmarks. While direct addition of $\tau_{A}$ to $\theta_{B}$ fails to transfer knowledge effectively, our method substantially closes the gap toward full fine-tuning, confirming its ability to identify and retain task-relevant directions. Notably, the relative improvement over naive transfer is even larger than in the vision setting, suggesting that sign-based filtering is particularly beneficial when source and target pre-training objectives differ more strongly. This is a challenging regime, since transferring updates between T5v1.1 and FLAN-T5 combines both pre-training and instruction-tuning mismatch. Despite this shift, GradFix remains effective with the same lightweight procedure used in vision. These results confirm that the method is architecture-agnostic in practice and supports reliable task-vector transport in the language domain as well.
5.2 Task Vector Transport for Model Merging
We evaluate GradFix in combination with model-merging methods, specifically Task Arithmetic (ilharco2023task) and TIES-Merging (yadav2023tiesmerging). We consider two settings: multi-task (one source model, multiple tasks) and multi-source (multiple source models, one task).
Multi-task experiments. For this setting, all task vectors are extracted from the same source pre-trained model $\theta_{A}$ (same pre-training, different downstream tasks) and transported to a fixed target pre-trained model $\theta_{B}$. For task vectors from distinct tasks, we compare two pipelines. Mask–then–Merge: transport each task vector with GradFix and then merge. Merge–then–Mask: first merge task vectors into $\tau_{\text{merged}}$, then transport the merged vector using a consensus mask computed by estimating per-parameter anti-gradient signs on $\theta_{B}$ for each task and selecting the most frequent sign at each coordinate. We report results for this setting in Tab. 3. Here, $\theta_{B}+\tau_{A,j}$ and $\theta_{B}+\delta^{A}_{j}$ denote single-task references evaluated on task $j$, respectively using naive addition and GradFix. Direct Task Arithmetic and TIES merging without GradFix perform near zero-shot, indicating strong cross-model misalignment, while Merge–then–Mask gives the best results. Masking each vector first can discard coordinates that would complement each other after merging; merging first preserves them and lets GradFix align one coherent update with the target loss geometry.
Table 3: Merging experiments on ViT-B/16 in multi-task and multi-source settings.
Pipeline EUROSAT SVHN GTSRB RESISC45 DTD AVG
$\theta_{B}$ zero-shot $49.41$ $50.58$ $48.29$ $67.98$ $55.96$ $54.44$
[…]
$\theta_{B}+\tau_{A,j}$ $49.58$ $50.84$ $49.31$ $67.87$ $56.27$ $54.77$
$\theta_{B}+\delta^{A}_{j}$ $65.07$ $70.19$ $\mathbf{64.33}$ $71.42$ $\mathbf{58.51}$ $65.90$
Task Arithmetic
    Baseline $49.31$ $50.99$ $48.73$ $68.05$ $56.54$ $54.73$
[]  Mask-then-Merge $55.90$ $71.56$ $59.65$ $71.40$ $57.66$ $63.23$
[]  Merge-then-Mask $65.37$ $72.10$ $59.55$ $71.16$ $57.07$ $65.05$
TIES-Merging
    Baseline $49.15$ $50.75$ $48.95$ $67.97$ $56.54$ $54.67$
[]  Mask-then-Merge $50.41$ $64.28$ $54.05$ $69.51$ $57.13$ $59.08$
[]  Merge-then-Mask $\mathbf{65.62}$ $\mathbf{72.42}$ $62.73$ $\mathbf{71.57}$ $57.77$ $\mathbf{66.02}$
[…]
Task Arithmetic
    Baseline $36.94$ $35.09$ $30.77$ $50.54$ $44.73$ $39.61$
[]  Merge-then-Mask $12.52$ $15.94$ $4.20$ $3.97$ $2.93$ $7.91$
[]  Mask-then-Merge $\mathbf{65.96}$ $\mathbf{72.97}$ $\mathbf{65.80}$ $71.30$ $61.01$ $\mathbf{67.41}$
TIES-Merging
    Baseline $12.69$ $10.39$ $2.85$ $5.81$ $16.33$ $9.61$
[]  Merge-then-Mask $14.46$ $15.94$ $3.09$ $3.35$ $2.66$ $7.90$
[]  Mask-then-Merge $65.17$ $72.58$ $65.36$ $\mathbf{71.40}$ $\mathbf{61.12}$ $67.13$
Multi-source experiments. In the multi-source setting, we use $K=5$ source models ${\{\theta_{A_{k}}\}}_{k=1}^{K}$ that are pre-trained on different data distributions and then fine-tuned on the same downstream task. We transport all resulting source task vectors to a single fixed target model $\theta_{B}$, whose pre-training remains unchanged during transport. Here, Merge–then–Mask is not expected to help: because all vectors correspond to the same task, the gradient-sign mask on $\theta_{B}$ is shared across sources, so consensus masking after merging is effectively equivalent to masking each source separately. We therefore use Mask–then–Merge: transport each source vector first, then merge the transported vectors. Table 3 shows that this recovers performance from the collapse of direct merging and also improves over single-source transport, suggesting that combining multiple transported updates provides a more robust descent direction.
@@ 5.3 Masking Strategies
Table 4: Performance of $\theta_{B}$ with oracle or estimated gradient signs under different mask strategies: sign agreement retains matching signs, sign forcing aligns all signs, magnitude-scaled uses the product of task and gradient magnitudes, and random assigns signs uniformly. Results averaged over seeds with $|\mathcal{D}_{s}^{c}|=1$ on CLIP ViT-B/16. The table includes $\theta_{B}$ zero-shot and $\theta_{B}$ fine-tune as lower and upper reference bounds.
@@
$\theta_{B}+\delta^{\star}$ sign agreement $95.06$ $87.06$ $82.92$ $92.04$ $71.44$ $85.71$
sign forcing $97.95$ $93.51$ $95.94$ $96.60$ $80.59$ $92.92$
magnitude-scaled $49.92$ $67.94$ $51.63$ $50.78$ $56.01$ $55.25$
$\theta_{B}+\delta^{A}$ sign agreement (ours) $61.94$ $70.05$ $60.89$ $71.07$ $58.32$ $64.45$
sign forcing $61.32$ $70.10$ $60.91$ $70.52$ $58.05$ $64.18$
magnitude-scaled $49.51$ $68.06$ $49.20$ $50.71$ $56.03$ $54.70$
$\theta_{B}+\delta^{A}$ random $49.49$ $67.97$ $48.41$ $50.54$ $56.06$ $54.50$
@@
To analyze the effect of different mask construction strategies on the transport of the task vector $\tau_{A}$, we compare our primary masking method (sign agreement), with three alternatives: sign forcing, magnitude-scaled, and random masks. Table 4 reports results using 1 sample per class on ViT-B/16, averaged across multiple random seeds.
@@
For both $\delta^{A}$ (Eq. 5) and $\delta^{\star}$ (Eq. 2), the mask $\bm{m}$} determines which directions of $\tau_{A}$ are retained. Here, $i$ indexes parameter coordinates. In sign agreement, $\bm{m}$} retains only the coordinates whose signs match those of the reference as in Eq. 4. In sign forcing, all signs of $\tau_{A}$ are aligned with the signs of the anti-gradient estimator $\hat{{\bm{s}}}$, obtaining:
$m_{i}^{sf}=\operatorname{sign}(\tau_{A,i})\cdot\hat{s}_{i},\qquad\delta_{i}^{sf}=\alpha\,(m_{i}^{sf}\tau_{A,i})=\alpha\,|\tau_{A,i}|\,\hat{s}_{i}.$ (11)
@@
This flips entries that disagree with the reference and applies a forced-sign update. When the oracle $\tau_{B}$ is used as the reference, sign forcing generally outperforms sign agreement, as fully leveraging the true task direction maximizes transfer. In contrast, using few-shot anti-gradient-based estimates from $\theta_{B}$, sign agreement performs slightly better than sign forcing. This is consistent with the fact that the anti-gradient estimate is noisy; forcing all directions can propagate errors, while keeping only agreeing entries provides a more reliable mask. We also evaluated a magnitude-scaled strategy to determine whether leveraging magnitude information offers additional benefits. This strategy computes the mask based on the magnitude of both the source task vector and the target reference vector $\rho$ (where $\rho=\tau_{B}$ for the oracle and $\rho=-g$ for the estimate):
$m_{i}^{ms}=\max(0,\tanh(\tau_{A,i}\cdot\rho_{i})),\qquad\delta_{i}^{ms}=\alpha\,(m_{i}^{ms}\tau_{A,i}).$ (12)
This approach assigns a mask value near $1.0$ only when components match in sign and possess significant magnitude, while suppressing mismatches. However, as shown in Tab. 4, this magnitude-aware strategy consistently underperforms sign agreement. Crucially, this failure persists even in the oracle setting ($\delta^{\star}$), which suggests that while directions are transferable across different pre-trainings, parameter magnitudes are highly specific to the local loss geometry of each basin. Consequently, enforcing magnitude consistency acts as an overly aggressive filter, discarding valid updates where the models agree on direction but differ on scale. In the limited-data regime ($\delta^{A}$), this issue is further exacerbated by the noise inherent in estimating gradient magnitudes from small datasets $\mathcal{D}_{s}$, confirming that the sign structure serves as the most robust proxy for alignment.
Finally, we evaluate random mask, where signs are sampled uniformly from $\mathcal{U}\{-1,+1\}$. Like magnitude-scaled masking, this approach performs close to the zero-shot baseline. This highlights that although sign-based masking outperforms magnitude-based filtering, the gain is not due to masking alone: random sign choices offer no useful signal. Thus, effective transfer relies strictly on the precise geometric alignment provided by the target anti-gradients, rather than just the sparsity of the update.
5.4 Sensitivity to the Scaling Coefficient
Figure 2: Accuracy across different $\alpha$ values for mean and majority-vote sign estimation.
We investigate the sensitivity of masked transport to the scaling factor $\alpha\in(0,1]$, providing a proxy for how compatible and robust the transported task vector is with the target backbone. In addition to our proposed majority voting strategy, we consider a baseline where the estimated sign is taken as the sign of the averaged anti-gradient (mean). Results are reported in Fig. 2.